OpenStudy (genny7):

The diameter of a circle is 14 yards. Find the area of the circle.

Nnesha (nnesha):

formula to find area of a circle \[\huge\rm A= \pi r^2\] where r =radius and radius is half of diameter r=d/2

OpenStudy (hyperlunatic):

Can I tell?

OpenStudy (hyperlunatic):

no I won't

Nnesha (nnesha):

sorry to say but no please let her solve it she is smart!

OpenStudy (hyperlunatic):

she should figure it out

Nnesha (nnesha):

right :=)

OpenStudy (genny7):

Okay got the answer. The answer i got is 153.86 yards squared

imqwerty (imqwerty):

can i tell the wrong answer (:

imqwerty (imqwerty):

aww :(

OpenStudy (hyperlunatic):

Why are so many people veiwing one question?

imqwerty (imqwerty):

cause of me

Nnesha (nnesha):

one question plz what is the radius ?? :=) last question :=)

OpenStudy (hyperlunatic):

no...

imqwerty (imqwerty):

cause of u

OpenStudy (genny7):

The radius would be 7

OpenStudy (hyperlunatic):

sshhh!!!

OpenStudy (anonymous):

that one is very simple

OpenStudy (anonymous):

its not hard

Nnesha (nnesha):

i was asking de person who posted the question. and yes genny that's right good work

OpenStudy (genny7):

Okay thanks so much @Nnesha :)

Nnesha (nnesha):

yw :=)